qwevery:

A zookeeper predicted the weight of a new baby elephant to be 259 pounds when it was born. The elephant actually weighed 291 pounds at birth. What was the percent error of the zookeeper's prediction? A. 32% B. 11.00% C. 12.36% D. 89.00%

Joe348:

wait do you have to round?

Caelum:

So you would multiply all, like so.; \[259*32~percent=?\] \[259*11~percent=?\] \[259*12.36~percent=?\] \[259*89~percent=?\]

qwevery:

tenths? idk-

Caelum:

We can obviously eliminate A.) and D.), they are too high.

Joe348:

just do 22 divided by 2

Joe348:

I got the explanation to below...

Caelum:

\[259*11~percent=28.49\] \[259+28.49=287.49\] Possibly, let's see the next one.

Joe348:

Joe348:

there you go dude

Caelum:

\[259*12.36~percent=32.0124\] \[32.0124+259=291\] Which means this one is correct.!

qwevery:

thank u:)

Caelum:

That is wrong @joe, and your welcome!

Caelum:

That doesn't even relate to the question.

Joe348:

no its not

Joe348:

your not even doing the percentage equation right

Caelum:

Omg, your screenshot doesn't even relate to the question...

Joe348:

okay you do you dude

AZ:

The numbers are different but that's the correct method to answer the question You won't always have answer choices \(\text{Percent error}= \dfrac{|\text{predicted - actual}|}{\text{actual}} \times 100\text{%}\) The predicted weight was 259 The actual weight was 291 pounds

Caelum:

If you look at your screenshot, look at the predicted vaule. Then, look at this question's predicted vaule.